A rectangle ABCD which measures 18 by 24 units is folded once, perpendicular to



diagonal AC, so that the opposite vertices A and C coincide. Find the length of the fold.




1
Expert's answer
2022-04-28T11:48:23-0400


ABCD - rectangle

AB = CD = 18

BC = AD = 24

MN perpendicular of diagonal AC

AO = OC = AC/2

MO = ON = x

From the Pythagorean theorem AC2 = AB2 + BC2

"AC = \sqrt{AB^2 + BC^2}\\
AC = \sqrt{18^2 + 24^2} = 30,\:then\:OC = 30/2 = 15"


From the triagle ABC

"\angle BCA = \theta\\
tan\theta = \frac{AB}{BC} = \frac{18}{24} = \frac{3}{4}"


From the triagle MCO

"\angle MCO = \theta\\
tan\theta = \frac{OM}{OC} = \frac{x}{15}, then\\
\frac{x}{15} = \frac{3}{4}\\
4x = 15\times3\\
x = 45/4 = 11.25,\: then\\
MN = 2\times x = 2\times11.25 = 22.5"

Answer:

Lenght of fold = 22.5